The speed of light in salt (sodium chloride) is 1.96 x10 8m/s . Calculate the index of refraction for sodium chloride. 



1
Expert's answer
2021-06-24T08:02:08-0400

Gives

Light velocity in medium

"v" ="1.96\times10^8m/sec"

Light velocity in vacuum

"c=3\times10^8m/sec"

n=?

We know that


"n=\frac{c}{v}"

Put value

"n=\frac{3\times10^8}{1.96\times10^8}=1.54"